Tate bought a compact disc for $17.99. The sales tax on the disc was t dollars. What was the total cost including sales tax? Can you show work and explain it please. Thank you
Answer by DrBeeee(684)

The price (cost) of the disc is $17.99.
Normally you are given a sales tax rate, say 9%. Then you take 9% of the sale price of $17.99 and get
(1) tax = 0.09*17.99 or
(2) tax = 1.6191 or to the nearest penny
(3) tax = 1.62
Then you add the tax to the price of the disc to get your total cost
(4) cost = 17.99 + 1.62 or the cost is $19.61 - pay the clerk!
In your problem you a simply told the amount of the tax which is the same as our (3) above. It is not a number but a variable, t. So to get the total cost in your problem add the tax, t, to the price $17.99 and get your total cost of
(5) Cost = 17.99 + t
Answer: the total cost, including the tax, is $17.99 plus t.